OpenStudy (anonymous):

the perimeter of a rectangle is 108 m. the length is 6 m more than twice the width. find the dimensions

OpenStudy (anonymous):

P=108 m Let w is the width and 2w+6 be the length. P=2l+2w 108=2(2w+6)+2w 108=4w+12+2w 108-12=6w 96=6w w=16 m 2w+6=length length=2(16)+6=32+6=38 m